# Demonstrating Wheeled Lab: Modern Sim2Real for Low-cost, Open-source Wheeled Robotics

### v1

@@ Abstract
Simulation has been pivotal in recent robotics milestones and is poised to play a prominent role in the field’s future. However, recent robotic advances often rely on expensive and high-maintenance platforms, limiting access to broader robotics audiences. This work introduces Wheeled Lab, a framework for the low-cost, open-source wheeled platforms that are already widely established in education and research. Through integration with Isaac Lab, Wheeled Lab introduces modern techniques in Sim2Real, such as domain randomization, sensor simulation, and end-to-end learning, to new user communities. To kickstart education and demonstrate the framework’s capabilities, we develop three state-of-the-art policies for small-scale RC cars: controlled drifting, elevation traversal, and visual navigation, each trained in simulation and deployed in the real world. By bridging the gap between advanced Sim2Real methods and affordable, available robotics, Wheeled Lab aims to democratize access to cutting-edge tools, fostering innovation and education in a broader robotics context. The full stack, from hardware to software, is low cost and open-source.
I Introduction
@@
The robotics community has made remarkable strides in recent years, boasting advances that enable quadrupeds to hike mountain trails and perform parkour [35, 26], autonomous drones to race at champion-level ability [31], and robot hands to solve Rubik’s Cubes [37]. Shared across these feats is the process of simulated policy learning and direct deployment in the real world, termed Sim2Real.
@@
However, these remarkable Sim2Real innovations now target expensive, space-intensive, and high-maintenance systems, including quadrupeds, humanoids, drones, and dexterous manipulators. Such pricey, high-end platforms make it impossible for the general population of robotics builders and users to access the state-of-the-art techniques they employ, such as domain randomization and sensor simulation [31, 28, 33, 26].
@@
In contrast, small-scale autonomous RC cars offer low-cost [^1] About 3000 USD or less. See [41] for a comparative analysis. yet performative platforms that can leverage modern robotics methods, as we demonstate in this paper. Used by classrooms, national racing competitions, and enthusiasts, platforms like the F1Tenth [38] have established themselves as go-to, entry-level platforms for beginning roboticists and researchers alike. Crucially, these platforms still provide sufficient onboard real estate to house components like the battery and compute necessary for modern algorithms as well as sufficient mobility to exercise complex dynamics and environment interactions, like drifting or off-roading [48, 24, 25, 14]. In short, small-scale wheeled robots have high algorithmic potential at relatively low cost and hardware complexity.
@@
A survey of existing ecosystems for low-cost wheeled platforms indicates that current simulation capabilities are isolated and outdated (see Table I). As a result, education [27] and research opportunities [19] limit practitioners to model-based or model-free methods with low-fidelity simulation. With the advent of high-fidelity, open-source simulators like Isaac Lab [36], which are geared toward modern robotic learning, these limited ecosystems no longer need to be the status quo.
@@
This work aims to support education and research by making modern robotics methods available to the broader community. To accomplish this, we contribute a Sim2Real framework for open-source wheeled robots that has been integrated with a state-of-the-art, research-grounded simulation ecosystem, Isaac Lab. Modern methods we implement and evaluate include: massive parallelization, domain randomization, sensor simulation, and end-to-end learning. We demonstrate these methods using three policy types trained in simulation and deployed on low-cost, open-source platforms:
@@
    The Elevation Policy ($\boldsymbol{\pi}_{\text{elev}}$) traverses three-dimensional features with spatial reasoning. This demonstration highlights end-to-end training for modern tasks that tightly couples perception and control. Wheeled Lab is the first to provide integration of an accessible elevation-based Sim2Real pipeline with massive parallelization.
@@
    The Visual Policy ($\boldsymbol{\pi}_{\text{vis}}$) traverses visual features with semantic understanding. This demonstration highlights simulation, end-to-end training, and deployment with cameras, a sensor modality gaining substantial research attention due to its pre-training potential. Wheeled Lab provides a lower cost alternative for visual Sim2Real compared to existing integrations.
Demonstrated policies also exhibit generalization to robustness and recovery behaviors during testing.
@@
II Intended Demonstration
We will run the policies on three separate platforms in front of the live audience. Where necessary, our evaluations of each policy use simple household objects and can also be brought to the demonstration. Running onboard and in real-time, each task and policy can also demonstrate their robustness and recovery behaviors. We will invite the audience to (safely) perturb the scene and robot during the demonstration. In particular, the drifting task can be deceptively challenging. The live audience will be invited to attempt the task themselves to experience the task’s complexity.
Importantly, these platforms are still entirely low-cost and open-source. We will invite the audience to experience the simple construction and design to help convey its low maintenance costs for use in education and research.
If time permits, the audience may also train their own models using the ecosystem and deploy them on our platforms.
III Related Work
@@
Wheeled Lab 3D Depth, RGB 16 DOF + Collisions High RSL, SB3 HOUND, MuSHR, F1Tenth
@@
CARLA [17] 3D Depth, RGB 16 DOF + Collisions Steering Only SB3
@@
TABLE I: Comparisons of existing ecosystems on their capabilities. Various simulation, learning, and deployment ecosystems have been integrated with accessible wheeled platforms for Sim2Real. However, these ecosystems are noticeably isolated from the research community and, where applicable, have outdated features.
@@
Considerable prior work in Sim2Real has addressed open-source wheeled robots, including [14, 39, 6, 40, 49, 13, 48, 18]. Some of these works produce one of $\boldsymbol{\pi}_{\text{drift}}$, $\boldsymbol{\pi}_{\text{elev}}$, or $\boldsymbol{\pi}_{\text{vis}}$ with partial or full Sim2Real. While the individual accomplishments of these works motivate our choice of demonstrations, they also help us highlight that varied state-of-the-art approaches can be coherently integrated under the framework presented here.
Drifting Policy. Drifting is a challenging task in both optimal control and reinforcement learning (RL) [22, 9, 16]. It is a fundamentally unstable maneuver whose control solutions are traditionally determined by vehicle and ground parameters [11]. In many cases (ours included), researchers are unable to demonstrate a controlled drift for imitation-based methods [13].
@@
These challenges make drifting an appealing task for RL. RL-based approaches have thus far been achievable only through further real-world fine-tuning [13, 48]. In this work, we show that aggressive domain randomization, perturbation simulation, and massive parallelization can enable direct transfer of a drifting policy. To reduce the barrier-to-entry, we deliberately distance our methods from techniques such as gain tuning and extensive system-identification that require additional domain expertise and tooling.
@@
Elevation Policy. It has become standard practice for locomotion policies on outdoor terrain to be trained with elevation maps [26]. In parallel, autonomous off-road vehicles have also demonstrated the importance of incorporating elevation in model-based methods [25, 20, 21]. Generally, these elevation-based tasks address challenges that arise at the interface of agent morphology and perception, which are often difficult to model. Therefore, lower-cost wheeled platforms have traditionally been oriented toward flat terrain [38, 6], revealing a gap in methodology between the broader community and the state-of-the-art. Notable exceptions include [14, 44, 50]. However, no work before this integrates these platforms with RL-facing simulators.
@@
Visual Policy. Improvements in camera and photorealistic scene simulation have advanced policies capable of semantic reasoning, bridging the gap between pre-trained visual models and embodied actions [30, 53]. Demonstrations of small-scale platforms with pre-trained visual models have also been investigated [44]. In general, cameras provide an approachable and interpretable sensor modality for modern robotics education and research. Unfortunately, similar to elevation policies, visual policy pipelines on low-cost mobile platforms are not readily available even though the success of proprietary camera-based Sim2Real platforms suggests a clear demand [6]. This work demonstrates simulation, end-to-end training, and deployment with cameras.
@@
Isaac Lab Simulation Framework. This work builds on Isaac Lab, an open-source, widely adopted, and rapidly growing simulation framework for robotics research [36, 33] that is heavily supported by industry and academia alike [47, 1, 2, 4, 3, 33]. Isaac Lab has become a primary learning framework of choice for recent robotics research due to its impressive performance and ease-of-use features [34, 26, 46, 51].
@@
Figure 1: Modular training framework imagined as the assembly of a puzzle. Three main components make up training: Run, Agent, Environment. Outlined puzzle pieces (e.g. Observation, Reward) are sub-components whose behaviors are defined by the pieces below them. For instance, High Speed is a reward setting in all of our RL environments, as denoted by the star (Refer to caption) icon. Wheel (Refer to caption), camera (Refer to caption), and mountain (Refer to caption) icons indicate specific design choices for $\boldsymbol{\pi}_{\text{drift}}$, $\boldsymbol{\pi}_{\text{vis}}$, and $\boldsymbol{\pi}_{\text{elev}}$, respectively. Components shown are non-exhaustive.
IV Wheeled Lab
Wheeled Lab bridges low-cost wheeled platforms with Isaac Lab (see LABEL:fig:fig1). To integrate with Isaac Lab, we simply define a hierarchy of abstractions. Some of the highest-level abstractions are illustrated in Figure 1 (e.g. Scene, Observation, Reward). Through iterative testing of the Sim2Real pipelines in this work, we have developed reliable baseline configurations of these abstractions which are amenable specifically to low-cost wheeled platforms. These definitions make up the platforms’ training environments in Isaac Lab.
In addition to training, Wheeled Lab itself further provides scripts for the evaluation of the pipeline, such as visualization, benchmarks, and reporting. These components are important aspects of the Sim2Real process which helps practitioners diagnose and iterate on algorithmic problems. For example, visualization of the value network on real data can help inform reward shaping and observation design [44].
To integrate with a platform, trained policies interface through ROS. We provide an example implementation of a node that builds the required observations from relevant topics and passes them forward to the trained network.
V Demonstrations
Figure 2: Family of HOUND platforms used for a robotics class at the University of Washington.
@@ V Demonstrations
For the demonstrations in this work, we use the HOUND [45] for $\boldsymbol{\pi}_{\text{drift}}$ and the MuSHR [^2] This work’s MuSHR uses the same casing as the HOUND but is otherwise a MuSHR in terms of mechanics and electronics. [43] for $\boldsymbol{\pi}_{\text{elev}}$ and $\boldsymbol{\pi}_{\text{vis}}$. The HOUND’s estimated total price is about 3000 USD for a complete autonomy package with sensing (i.e., LiDAR, RGB) and compute (i.e., Jetson Orin NX). The MuSHR is estimated at 930 USD. Within the context of our work, the HOUND platform’s hardware is nearly equivalent to that of the F1tenth platform (3800 USD). A single NVIDIA RTX 3080 GPU is used for training.
To demonstrate the capabilities of the framework, Wheeled Lab implements, deploys, and evaluates three state-of-the-art policy types: $\boldsymbol{\pi}_{\text{drift}}$, $\boldsymbol{\pi}_{\text{elev}}$, and $\boldsymbol{\pi}_{\text{vis}}$ on the HOUND platform.
In each task, we compare the policy trained with modern Sim2Real techniques (e.g. domain randomization, perturbations, etc) against a baseline policy trained without them to demonstrate their effectiveness in Sim2Real transfer in the presented domains. Settings for drifting and elevation tasks can be found in Table II. All other configuration settings not mentioned (e.g. rewards, actuator parameters, articulations) can be assumed to be the same between $\boldsymbol{\pi}$ and $\overline{\boldsymbol{\pi}}$. Each baseline is a liberal representation of what training ecosystems currently provide to low-cost wheeled platforms as shown in Table I. We use $\overline{(\cdot)}$ to denote the baseline version of a policy (e.g. $\overline{\boldsymbol{\pi}}_{\text{drift}}$). Across all policies, we use Proximal Policy Optimization (PPO) [42] as implemented in the RSL RL library to train the agent in simulation [32].
@@
TABLE II: Training Settings. Baseline policies $\boldsymbol{\pi}_{\text{drift}}$ and $\boldsymbol{\pi}_{\text{elev}}$ (liberally) reflect capabilities currently available to low-cost open-source wheeled platforms.
V-A Drifting
Figure 3: A control strategy arises for drifting through turns. Left column: (Overlay: colored trajectories over multiple runs.) One run is spotlighted for this visualization. Arrows indicate the direction of travel, and colorbar denotes speeds (in m/s). Middle column: Magnification of drifted turns with markers visualizing vehicle orientation, steering (front wheel direction), and throttle (rear wheels; blue: $-1$, pink: $+1$). Right column: Slip angle, speed, steering, and throttle are plotted for each turn. To initiate the drift, the platform quickly cuts the throttle to destabilize the rear wheels then steers inwards sharply to throw them outwards. With the platform now facing the track center, it throttles through the remainder of the turn while counter-steering to control its residual angular momentum from the initial maneuver. The entire sequence occurs in just over 1 second. Visualization inspired by [16].
Figure 4: Captured trajectories from the baseline drift policy in real experiments. Baseline is unable to complete a valid turn without crashing or spinning out.
Background. Drifting is characterized by a vehicle’s side-slip angle, the angle between the heading and the velocity, as it maneuvers through a turn. In high-performance driving, drifting allows the vehicle to take much sharper turns at higher speeds than otherwise allowable by the no-slip turn radius [5]. The total mass, center-of-mass, friction properties of the surface and the wheels, actuator properties, state estimation, steering linkages, and suspension are all relevant factors in the narrow dynamics of a controlled drift.
The sensitivity of these dynamics becomes especially obvious when we, the researchers, attempt the task manually despite the deceptively simple action space (see website for videos). When the maneuver is not executed correctly, the vehicle spins uncontrollably off track, losing traction and usually requiring a reset (also known as a spin out).
Setup. State estimation is provided primarily by onboard VIO. A motion capture system is used for evaluation and occasional (1 hz) integrator-drift correction due to space limitations of our indoor testing environment. Successful VIO-only runs not used for data collection can be found on our website. To enable the platform to be physically capable of drifting, tape is wrapped around the platform’s rubber wheels to reduce friction and the base is converted to a rear-wheel drive.
@@ V-A Drifting
Results. We find that the baseline is unable to make a stable turn, much less complete a full lap (Figure 4). However, it occasionally attempts to counter-steer despite showing no indication during training. Baseline training runs consistently do not discover the drifting mode. Motor cogging, noisy state estimates, constant slipping, and steering biases prove to cause significant covariate shift.
@@
On the other hand, $\boldsymbol{\pi}_{\text{drift}}$ is able to complete full laps. In addition, we observe evidence of high robustness in $\boldsymbol{\pi}_{\text{drift}}$ not yet seen in previous literature [48, 16]. Namely, when the platform does spin out, instead of attempting to regain control by cutting the throttle, the policy maintains its angular momentum for a full spin (or more) before returning to the track. A visualization of the results gives insight into the policy’s precise control in Figure 3. The maximum (controlled) slip angle is $58^{\circ}$ and the average speed is about $1.6$ m/s.
Further details about the design and implementation of $\boldsymbol{\pi}_{\text{drift}}$ can be found in Appendix B-A.
V-B Elevation
Figure 5: Captured trajectories of elevation policy in real experiments. Top: $\boldsymbol{\pi}_{\text{elev}}$ demonstrates spatial reasoning through ramp traversal and obstacle avoidance in one sequence using only a local elevation map. Bottom: Nearing a rollover, the Baseline $\overline{\boldsymbol{\pi}}_{\text{elev}}$ falls off the ramp and then crashes into it.

### v2

# Wheeled Lab: Modern Sim2Real for Low-cost, Open-source Wheeled Robotics

@@ Abstract
Reinforcement Learning (RL) has been pivotal in recent robotics milestones and is poised to play a prominent role in the future. However, these advances can rely on proprietary simulators, expensive hardware, and a daunting range of tools and skills. As a result, broader communities are disconnecting from the state-of-the-art; education curricula are poorly equipped to teach indispensable modern robotics skills involving hardware, deployment, and iterative development. To address this gap between the broader and scientific communities, we contribute Wheeled Lab, an ecosystem which integrates accessible, open-source wheeled robots with Isaac Lab, an open-source robot learning and simulation framework, that is widely adopted in the state-of-the-art. To kickstart research and education, this work demonstrates three state-of-the-art zero-shot policies for small-scale RC cars developed through Wheeled Lab: controlled drifting, elevation traversal, and visual navigation. The full stack, from hardware to software, is low-cost and open-source. Videos and additional materials can be found at: https://uwrobotlearning.github.io/WheeledLab/
11footnotetext: Corresponding author: than123@uw.edu
1 Introduction
@@
The robotics community has made remarkable strides in recent years, boasting advances that enable quadrupeds to hike mountain trails and perform parkour [5, 6], autonomous drones to race at champion-level ability [7], and robot hands to solve Rubik’s Cubes [8]. Shared across these feats is simulated policy learning followed by direct deployment in the real world, termed Sim2Real.
@@
However, these remarkable Sim2Real innovations currently target expensive, space-intensive, and high-maintenance systems, including quadrupeds, humanoids, drones, and dexterous manipulators. Such platforms make it difficult for the general population of robotics builders and users to access the state-of-the-art techniques they employ, such as domain randomization and sensor simulation [7, 9, 10, 6].
@@
In contrast, small-scale autonomous RC cars offer low-cost [^1] About 3000 USD or less. See [11] for a comparative analysis. yet performative platforms that can leverage modern robotics methods, as we demonstrate in this paper. Used by classrooms, national racing competitions, and enthusiasts, platforms like F1Tenth [3] have established themselves as go-to, entry-level platforms for beginning roboticists and researchers alike. Crucially, these platforms still provide sufficient onboard real estate to house components like the battery and compute necessary for modern algorithms, as well as sufficient mobility to exercise complex dynamics and environment interactions like drifting or off-roading [12, 13, 14, 15]. In short, small-scale wheeled robots have high algorithmic potential at relatively low cost and hardware complexity.
@@
A survey of existing ecosystems for low-cost wheeled platforms indicates that current simulation capabilities are isolated and outdated (see Table 1). As a result, education [16] and research opportunities [17] currently limit practitioners to model-based or model-free methods with low-fidelity simulation. With the advent of high-fidelity, open-source simulators like Isaac Lab [4], which are geared toward modern robotic learning, these limited ecosystems no longer need to be the status quo.
@@
This work aims to support education and research by making modern robotics methods available to the broader community. To accomplish this, we contribute a Sim2Real framework for open-source wheeled robots that we have integrated with a state-of-the-art, research-grounded simulation ecosystem, Isaac Lab. Modern methods we implement and evaluate include: massive parallelization, domain randomization, sensor simulation, and end-to-end learning. We demonstrate these methods using three policy types trained in simulation and deployed on low-cost, open-source platforms:
@@
    The Elevation Policy ($\boldsymbol{\pi}_{\text{elev}}$) traverses three-dimensional features with spatial reasoning. This demonstration highlights end-to-end training for modern tasks that tightly couple perception and control. Wheeled Lab is the first to provide integration of an accessible elevation-based Sim2Real pipeline with massive parallelization.
@@
    The Visual Policy ($\boldsymbol{\pi}_{\text{vis}}$) traverses visual features with semantic understanding. This demonstration highlights simulation, end-to-end training, and deployment with cameras, a sensor modality gaining substantial research attention due to its pre-training potential. Wheeled Lab provides a lower-cost alternative for visual Sim2Real compared to existing integrations.
@@
2 Related Work
@@
Wheeled Lab 3D Depth, RGB
16 DOF
+ Collisions
High RSL, SB3
HOUND, MuSHR,
F1Tenth
@@ 16 DOF
CARLA [18] 3D Depth, RGB
16 DOF
+ Collisions
Steering
Only
SB3
@@
Table 1: Comparisons of existing ecosystems on their capabilities. Various simulation, learning, and deployment ecosystems have been integrated with accessible wheeled platforms for Sim2Real. However, these ecosystems are noticeably isolated from the research community and, where implemented, have outdated features.
@@
Considerable prior work in Sim2Real has addressed open-source wheeled robots, including [15, 24, 23, 25, 26, 27, 12, 28]. Some of these works produce one of $\boldsymbol{\pi}_{\text{drift}}$, $\boldsymbol{\pi}_{\text{elev}}$, or $\boldsymbol{\pi}_{\text{vis}}$ with partial or full Sim2Real. The individual accomplishments of this research both motivate our choice of demonstrations and help us highlight that varied state-of-the-art approaches can be coherently integrated under the framework presented here.
@@
Drifting Policy. Drifting is a challenging task in both optimal control and reinforcement learning (RL) [29, 30, 31]. It is a fundamentally unstable maneuver whose control solutions are traditionally determined by vehicle and ground parameters [32]. In many cases (ours included), researchers are unable to demonstrate a controlled drift for imitation-based methods [27]. These challenges make drifting an appealing task for RL. RL-based approaches have thus far been achievable only through further real-world fine-tuning [27, 12]. In this work, we show that aggressive domain randomization, perturbation simulation, and massive parallelization can enable direct transfer of a drifting policy. To reduce the barriers-to-entry, we deliberately distance our methods from techniques such as gain tuning and extensive system identification that require additional domain expertise and tooling.
@@
Elevation Policy. It has become standard practice for locomotion policies on outdoor terrain to be trained with elevation maps [6]. In parallel, autonomous off-road vehicles have also demonstrated the importance of incorporating elevation in model-based methods [14, 33, 34]. Generally, these elevation-based tasks address challenges that arise at the interface of agent morphology and perception, which are often difficult to model. Therefore, lower-cost wheeled platforms have traditionally been oriented toward flat terrain [3, 23], revealing a gap in methodology between the broader community and the state-of-the-art. Notable exceptions include [15, 35, 36] but do not offer as extensive community support for RL-based approaches.
@@
Visual Policy. Improvements in camera and photorealistic scene simulation have advanced policies capable of semantic reasoning, bridging the gap between pre-trained visual models and embodied actions [37, 38]. Demonstrations of small-scale platforms with pre-trained visual models have also been investigated [35]. In general, cameras provide an approachable and interpretable sensor modality for modern robotics education and research. Unfortunately, similar to elevation policies, visual policy pipelines on low-cost mobile platforms are not readily available even though the success of proprietary camera-based Sim2Real platforms suggests a clear demand [23]. This work demonstrates simulation, end-to-end training, and deployment with cameras on a low-cost platform.
@@
Isaac Lab Simulation Framework. This work builds on Isaac Lab, an open-source, widely adopted, and rapidly growing simulation framework for robotics research [4, 10] that is heavily supported by industry and academia alike [39, 40, 41, 42, 43, 10]. Isaac Lab has become a primary learning framework of choice for recent robotics research due to its impressive performance and ease-of-use features [44, 6, 45, 46, 47].
@@
Figure 1: Modular training framework imagined as the assembly of a puzzle. Three main components comprise training: Run, Agent, Environment. Outlined puzzle pieces (e.g., Observation, Reward) are sub-components whose behaviors are defined by the pieces below them. For instance, High Speed is a reward setting in all of our RL environments, as denoted by the star (Refer to caption) icon. Wheel (Refer to caption), camera (Refer to caption), and mountain (Refer to caption) icons indicate specific design choices for $\boldsymbol{\pi}_{\text{drift}}$, $\boldsymbol{\pi}_{\text{vis}}$, and $\boldsymbol{\pi}_{\text{elev}}$, respectively. Components shown are non-exhaustive.
3 Implementation
Wheeled Lab bridges low-cost wheeled platforms with Isaac Lab (see LABEL:fig:fig1). We extend Isaac Lab by designing infrastructure and assets which enforce reproducible experimentation. Some of the highest-level abstractions towards this effort are illustrated in Figure 1 (e.g., Scene, Observation, Reward), highlighting the inclusion of run parameters, architecture, and hyperparameters in addition to existing environment configurations. This design helps ensure reproducible, accessible, and portable experimentation settings and results, which is critical for practitioners to build upon existing work.
This work presents reliable and empirically validated baseline configurations of these abstractions, which are amenable specifically to low-cost wheeled platforms. As Sim2Real remains an iterative process, especially in low-cost settings, our implementations consider accessibility factors throughout the entire development process: limited onboard compute and sensing, visual diversity, system identification feasibility, access to evaluation terrain, access to open-source packages, and access to training resources. Further implementation details for each task can be found in Appendix B.
4 Experiments
Figure 2: Several autonomy platforms shared among students in a robotics class.
While we contribute the novelties in applied RL as discussed in Section 1, our experiments aim to highlight the disparity in open-source methodologies between the broader and scientific communities. Thus, we compare policies trained through Wheeled Lab to baselines which reflect what training ecosystems currently provide to low-cost wheeled platforms, as shown in Table 1. This includes ablations over techniques such as domain randomization, perturbations, parallelization capacity, and architectures. $\overline{(\cdot)}$ denotes the baseline version of a policy (e.g., $\overline{\boldsymbol{\pi}}_{\text{drift}}$). Across all policies, we use Proximal Policy Optimization (PPO) [48] as implemented in the RSL RL library to train the agent in simulation [49].
We implement, deploy, and evaluate three state-of-the-art policy types: $\boldsymbol{\pi}_{\text{drift}}$, $\boldsymbol{\pi}_{\text{elev}}$, and $\boldsymbol{\pi}_{\text{vis}}$ on the HOUND and MuSHR platforms. Settings for drifting and elevation tasks are shown in Table 2. All other configuration settings not mentioned (e.g., rewards, actuator parameters, articulations) can be assumed to be the same between $\boldsymbol{\pi}$ and $\overline{\boldsymbol{\pi}}$.
@@ 4 Experiments
Our experiments employ the HOUND [1] for $\boldsymbol{\pi}_{\text{drift}}$ and the MuSHR [2] for $\boldsymbol{\pi}_{\text{elev}}$ and $\boldsymbol{\pi}_{\text{vis}}$. The HOUND’s estimated total price is about 3000 USD for a complete autonomy package with sensing (i.e., LiDAR, RGB) and compute (i.e., Jetson Orin NX). The MuSHR is estimated at 930 USD. A single NVIDIA RTX 3080 GPU is used for training. Isaac Lab also offers cloud-based training for broader audiences without personal GPU access.
Figure 3: A control strategy arises for drifting through turns. Left: (Overlay: colored trajectories over multiple runs.) One run is spotlighted for this visualization. Arrows indicate the direction of travel, and the colorbar denotes speeds (in m/s). Middle: Magnification of drifted turns with markers visualizing vehicle orientation, steering (front wheel direction), and throttle (rear wheels; blue: 1-1, pink: +1+1). Right: Slip angle, speed, steering, and throttle are plotted for each turn. To initiate the drift, the platform quickly cuts the throttle to destabilize the rear wheels and then steers inwards sharply to throw them outwards. With the platform now facing the track center, it throttles through the remainder of the turn while counter-steering to control its residual angular momentum from the initial maneuver. The entire sequence occurs in just over 1 second. Visualization inspired by [31].
Figure 4: Captured trajectories from the baseline drift policy in real experiments. The baseline is unable to complete a valid turn without crashing or spinning out.
4.1 Drifting
Background. Drifting is characterized by a vehicle’s side-slip angle, i.e., the angle between the heading and the velocity, as it maneuvers through a turn. In high-performance driving, drifting allows the vehicle to take much sharper turns at higher speeds than would be otherwise allowable by the no-slip turn radius [50]. The total mass, center-of-mass, friction properties of the surface and the wheels, actuator properties, state estimation, steering linkages, and suspension are all relevant factors in the narrow dynamics of a controlled drift.
The sensitivity of these dynamics becomes especially obvious when we (researchers) attempt the task manually despite the deceptively simple action space (see website for videos). When the maneuver is not executed correctly, the vehicle spins uncontrollably off track, losing traction (also known as a spin out) and usually requiring a reset.
Setup. State estimation is provided primarily by onboard Visual-Inertial-Odometry (VIO). We use a motion capture system for evaluation and occasional (1 hz) integrator-drift correction due to space limitations of our indoor testing environment. Successful VIO-only runs not used for data collection are shown on our website. To enable the platform to be physically capable of drifting, tape is wrapped around the platform’s rubber wheels to reduce friction and converted the base to a rear-wheel drive.
@@ 4.1 Drifting
Table 2: Training settings. Baseline policies $\boldsymbol{\pi}_{\text{drift}}$ and $\boldsymbol{\pi}_{\text{elev}}$ (liberally) reflect capabilities currently available to low-cost, open-source wheeled platforms.
@@
Results. We find that the baseline is unable to make a stable turn, much less complete a full lap (Figure 4). However, it occasionally attempts to counter-steer despite showing no indication during training. Baseline training runs consistently fail to discover the drifting mode. Motor cogging, noisy state estimates, constant slipping, and steering biases cause significant covariate shift.
@@
On the other hand, $\boldsymbol{\pi}_{\text{drift}}$ is able to complete full laps. In addition, we observe evidence of high robustness in $\boldsymbol{\pi}_{\text{drift}}$ not yet seen in previous literature [12, 31]. That is, when the platform does spin out, instead of attempting to regain control by cutting the throttle, the policy maintains its angular momentum for a full spin (or more) before returning to the track. A visualization of the results (Figure 3) gives insight into the policy’s precise control. The maximum (controlled) slip angle is $58^{\circ}$, and the average speed is about $1.6$ m/s. Appendix B.1 provides further details about the design and implementation of $\boldsymbol{\pi}_{\text{drift}}$.
4.2 Elevation
(a) Captured trajectories of elevation policy in real experiments.
(b) 3D view of global elevation map.
Figure 5: Comparison of policy $\boldsymbol{\pi}_{\text{elev}}$ and baseline $\overline{\boldsymbol{\pi}}_{\text{elev}}$ in simulation and real-world elevation tasks. Fig. 5(a) shows real-world execution: $\boldsymbol{\pi}_{\text{elev}}$(top) successfully avoids obstacles using local elevation, while $\overline{\boldsymbol{\pi}}_{\text{elev}}$ (bottom) avoids elevation features entirely. Fig. 5(b) visualizes the elevation map along with additional experiment result annotations. Trajectories from $\boldsymbol{\pi}_{\text{elev}}$ (yellow/blue) and $\overline{\boldsymbol{\pi}}_{\text{elev}}$ (purple/white) are overlaid. Policies only access their local $2.5\times 2.5$ m maps during evaluation.